OpenStudy (anonymous):

find derivative of ln (1/sqrtx)

OpenStudy (anonymous):

\[\ln 1/\sqrt{x}\]

OpenStudy (anonymous):

\[\ln (\frac{ 1 }{ \sqrt{x} })\]

OpenStudy (anonymous):

That right

OpenStudy (anonymous):

well there were no parentheses but yes thats what it looks like

OpenStudy (anonymous):

It just requires the chain rule, do you know of it?

OpenStudy (anonymous):

yes but getting hung up on this problem

OpenStudy (anonymous):

Can you show me what you did, so I can see where exactly you're having troubles?

OpenStudy (sidsiddhartha):

 this will help